monkey is decending from the branch of a tree with constant acceleration. If the breaking strength is 75% of the weight of the monkey, the minimum acceleration with which monkey can slide down without breaking the branch is (a) g (b) 3g/4 (c) g/4 (d) g/2How is last step a=g/4 plz explain in detail

Grade:11

1 Answers

RITWIK RAJ JHA
6 years ago
see here monkey is descending so ma will be acting downward so account.to free body diagram (which I can not draw here) mg-ma=3/4mg.so, solving it we get a=g/4 which is the maximum acceleration with which monkey can slide down without breaking the branch